# Constrained Denoising, Empirical Bayes, and Optimal Transport

### v1

@@ Abstract
Keywords: constrained Bayes estimation; $G$-modeling; Gaussian mixture model; error-in-variables regression; latent variable model; (smoothed) nonparametric maximum likelihood estimation; Wasserstein space
@@ 1.1 Problem Statement
More precisely, the goal of denoising is to construct a function $\delta:\mathbb{R}^{d}\to\mathbb{R}^{m}$ such that the risk $\mathbb{E}[\|\delta(Z)-\Theta\|^{2}]$ is small, where the expectation is taken over the joint distribution of $(\Theta,Z)$. If the distribution $G$ is known, then the best choice is the posterior mean $\delta_{\mathcal{B}}(z):=\mathbb{E}[\Theta\,|\,Z=z]$, called the Bayes denoiser (with respect to the squared error loss). If the distribution $G$ is not known, then it is often still possible to construct some function $\hat{\delta}_{\mathcal{B}}$, called an empirical Bayes (EB) denoiser (Robbins, 1956; Efron, 2019), which achieves a risk that is nearly equal to that of the Bayes denoiser.
@@
Similar considerations have been noted by several authors over the last hundred years, e.g., Eddington (1940); Louis (1984); Ghosh (1992); Loredo (2007); Bloom et al. (2017); Agarwal et al. (2020), and in textbooks [Raudenbush and Bryk, 2002, p. 88; Rao and Molina, 2015, p. 316].
@@ 1.2 Classical Denoising and Three Constrained Denoising Problems
We now describe the classical denoising problems and three constrained variants that are of interest in this work. First, we consider the (unconstrained) Bayes denoising problem:
$\begin{cases}\textnormal{minimize}&\mathbb{E}\left[\|\delta(Z)-\Theta\|^{2}\right]\\ \textnormal{over}&\delta:\mathbb{R}^{d}\to\mathbb{R}^{m}\end{cases}$ ($\mathcal{B}$)
whose solution is the posterior mean $\delta_{\mathcal{B}}(z)=\mathbb{E}[\Theta\,|\,Z=z]$. Next, we consider adding constraints to problem ($\mathcal{B}$).
@@
$\begin{cases}\textnormal{minimize}&\mathbb{E}\left[\|\delta(Z)-\Theta\|^{2}\right]\\ \textnormal{over}&\delta:\mathbb{R}^{d}\to\mathbb{R}^{m}\\ \textnormal{with}&\mathbb{E}[\delta(Z)]=\mathbb{E}[\Theta]\\ \textnormal{and}&\textnormal{Cov}(\delta(Z))=\textnormal{Cov}(\Theta).\end{cases}$ ($\mathcal{VCB}$)
Problem ($\mathcal{VCB}$) aims to find a denoiser $\delta$ that minimizes the Bayes risk subject to the constraint that the first two moments of $\delta$ match those of the latent variable $\Theta$, thus ensuring that the shrinkage in (1.2) is mitigated.
@@
$\begin{cases}\textnormal{minimize}&\mathbb{E}\left[\|\delta(Z)-\Theta\|^{2}\right]\\ \textnormal{over}&\delta:\mathbb{R}^{d}\to\mathbb{R}^{m}\\ \textnormal{with}&\delta(Z)\overset{\mathcal{D}}{=}\Theta.\end{cases}$ ($\mathcal{DCB}$)
@@
$\begin{cases}\textnormal{minimize}&\mathbb{E}\left[\|\delta(Z)-\Theta\|^{2}\right]\\ \textnormal{over}&\delta:\mathbb{R}^{d}\to\mathbb{R}^{m}\\ \textnormal{with}&\mathbb{E}[\psi_{\ell}(\delta(Z))]=\mathbb{E}[\psi_{\ell}(\Theta)]\quad\textnormal{ for all }1\leq\ell\leq k\end{cases}$ ($\mathcal{GCB}$)
where $\psi_{1},\ldots,\psi_{k}:\mathbb{R}^{m}\to\mathbb{R}$ is some prescribed set of measurable functions. This formulation is general enough to cover a wide variety of problems; for example, it encompasses ($\mathcal{VCB}$) when $m=1$ by taking $\psi_{1}(z):=z$ and $\psi_{2}(z):=z^{2}$, it also encompasses ($\mathcal{DCB}$) in a limiting sense when $\psi_{1},\psi_{2},\ldots$ are taken to be a suitable sequence of bounded measurable functions. The added generality is useful in some novel applications and substantially extends the scope of constrained denoising. (See Examples 3.93.10, and 3.11.)
@@ 1.3 Summary of Results
The main results of this paper concern the various constrained Bayes problems above, their EB approximations, some extensions, and some applications. We now describe these points in more detail.
@@
Distribution-Constrained Denoising. Next, we consider the distribution-constrained Bayes problem ($\mathcal{DCB}$). At the population level, we give conditions under which there is a uniquely defined distribution-constrained Bayes denoiser $\delta_{\mathcal{DCB}}$ (Theorem 3.8); at the empirical level, we provide an EB denoising scheme (Algorithm 2) which produces a denoiser $\hat{\delta}_{\mathcal{DCB}}$ that exhibits an explicit rate of convergence to the oracle constrained denoiser (Theorem 4.6). This work extends recent results in García-Trillos and Sen (2024) concerning the population-level problem; we provide rigorous statistical guarantees for an EB procedure which was stated (but not rigorously studied) in García-Trillos and Sen (2024, Appendix E). In Figure 1.1 (fourth column) we display the resulting denoisers in the simulation.
@@
General-Constrained Denoising. Third, we consider the general-constrained Bayes problem ($\mathcal{GCB}$), which encapsulates a wide variety of practical constraints (for instance, the fourth-moment constraints of Armstrong et al. (2022) or support constraints such as nonnegativity). At the population level, we give conditions under which there is a uniquely-defined general-constrained Bayes denoiser $\delta_{\mathcal{GCB}}$ (Theorem 3.12); at the empirical level, we provide an EB denoising scheme (Algorithm 3) which produces a denoiser $\hat{\delta}_{\mathcal{GCB}}$ that exhibits an explicit rate of convergence to the oracle constrained denoiser (Theorem 4.7).
@@
A fundamental flexibility of our paradigm is that it allows one to transform an arbitrary unconstrained EB denoiser into a suitable constrained EB denoiser, and the constrained denoiser inherits its rate of convergence from the unconstrained EB denoiser. For example, the unconstrained denoiser may arise via $G$-modeling, $F$-modeling, parametric EB, smoothed versions thereof, hybrids thereof, and more. This allows us to determine the rates of convergence for our constrained EB denoising procedures in many different settings of interest (e.g., light- and heavy-tailed $G$, Gaussian and Poisson likelihood $\{P_{\theta}\}_{\theta}$, conjugate parametric models), and we give a detailed analysis of the resulting rates of convergence in Appendix C. A notable example is the case of a nonparametric latent variable distribution and a Gaussian likelihood where we use nonparametric maximum likelihood (NPMLE) for $G$-modeling; our results imply that the rate of convergence of $\hat{\delta}_{\mathcal{VCB}}$ is nearly-parametric, and the rate of convergence of $\hat{\delta}_{\mathcal{DCB}}$ is slow (due to the slow minimax rate of nonparametric deconvolution, i.e., Carroll and Hall (1988); Zhang (1990); Fan (1991)).
@@
We also illustrate our methodology in two applications (Section 6). In an application to astronomy (Section 6.1), we use our results to denoise the distribution of latent chemical abundances in a large catalog of observed stars, furthering the analyses of Soloff et al. (2024); Ratcliffe et al. (2020). In an application to baseball (Section 6.2), we use our results to denoise the distribution of latent batting skill for players who are promoted from minor- to major-league play. Because these applications involve some degree of heterogeneity in the observations (e.g., heteroskedasticity in the Gaussian case) we include a detailed discussion (Section 5) of some considerations for extensions to the heterogeneous case.
@@ 1.4 Related Literature
Additionally, our results are related to some recent literature on distribution-constrained denoising, e.g., García-Trillos and Sen (2024). Indeed, our results provide rates of convergence for the distribution-constrained EB procedure described in García-Trillos and Sen (2024), which was motivated in part by recent literature on the “distortion-perception tradeoff” in computer vision (Blau and Michaeli, 2018; Freirich et al., 2021).
Our proposed methodology is closely tied to practical considerations in a number of application areas. In this paper, we specifically consider applications to astronomy (Section 6.1) and baseball (Section 6.2). Regarding astronomy, Loredo (2007) has argued that distributional constraints on denoising procedures are a fundamental feature of astrostatistics. Another application is to baseball statistics; while our specific application is novel, it complements a long tradition of EB methods in baseball (Efron and Morris, 1975; Brown, 2008; Jiang and Zhang, 2010; Gu and Koenker, 2017). We believe our proposed methodology will also be useful in several additional application areas, for example in the analysis of multisite experiments that are increasingly common in education research (Raudenbush and Bloom, 2015; Bloom et al., 2017; Lee et al., 2024), small-area estimation (Devine et al., 1994; Ghosh and Rao, 1994; Leyland and Davies, 2005) and error-in-variables regression (Carroll and Stefanski, 1990; Freedman et al., 2004; Bay, 1997; Thomas et al., 2011).
2 Notation and Preliminaries
In this section we review some preliminaries that will be used in our main results. For notation, let us fix $m\geq 1$ and $d\geq 1$, and take $G$ to be an unknown probability measure on $\mathbb{R}^{m}$ and $\{P_{\theta}\}_{\theta}$ to be a known family of probability measures on $\mathbb{R}^{d}$. Then we define the probability measure $F$ on $\mathbb{R}^{d}$ via $F(\textnormal{d}z)=\int_{\mathbb{R}^{m}}P_{\theta}(\textnormal{d}z)\textnormal{d}G(\theta)$. We write $(\Theta,Z)$ for a pair sampled from model (1.1), and $(\Theta_{1},Z_{1}),\ldots,(\Theta_{n},Z_{n})$ for independent, identically-distributed (i.i.d.) sequence of such pairs.
2.1 Bayes Denoisers and Risks
Problem ($\mathcal{B}$) has a unique solution $\delta_{\mathcal{B}}:\mathbb{R}^{d}\to\mathbb{R}^{m}$ given by the posterior mean $\delta_{\mathcal{B}}(z)=\mathbb{E}[\Theta\,|\,Z=z]$, and we denote its minimum value by $R_{\mathcal{B}}$; we refer to these as the (unconstrained) Bayes denoiser and the (unconstrained) Bayes risk, respectively. When we wish to emphasize the dependence on the latent variable distribution $G$, we will augment the notation and write $\delta_{\mathcal{B},G}$ and $R_{\mathcal{B},G}$, respectively.
We similarly write $\delta_{\mathcal{VCB}},\delta_{\mathcal{DCB}}$, and $\delta_{\mathcal{GCB}}$ for solutions to problems ($\mathcal{VCB}$), ($\mathcal{DCB}$), and ($\mathcal{GCB}$), respectively (when they exist), and $R_{\mathcal{VCB}},R_{\mathcal{DCB}}$, and $R_{\mathcal{GCB}}$ for minimal values of the corresponding problems.
2.2 Optimal Transport
@@
$\begin{cases}\textnormal{minimize}&\int_{\mathbb{R}^{k}}\|x-T(x)\|^{2}\textnormal{d}\mu(x)\\ \textnormal{over}&T:\mathbb{R}^{k}\to\mathbb{R}^{k}\\ \textnormal{with}&T_{\#}\mu=\mu^{\prime}.\end{cases}$ (2.1)
@@
$\begin{cases}\textnormal{minimize}&\int_{\mathbb{R}^{k}}\|x-x^{\prime}\|^{2}\textnormal{d}\pi(x,x^{\prime})\\ \textnormal{over}&\pi\in\Gamma(\mu;\mu^{\prime}).\end{cases}$ (2.2)
@@
In practice, a common approach to solving (2.1) is as follows. First, find a solution $\pi$ to (2.2). Second, define the function $\delta:\mathbb{R}^{k}\to\mathbb{R}^{k}$ via
$\delta(x^{\prime}):=\frac{\int_{\mathbb{R}^{k}}x\,\pi(\textnormal{d}x,x^{\prime})}{\int_{\mathbb{R}^{k}}\pi(\textnormal{d}x,x^{\prime})}$
called the barycentric projection of $\pi$; note that this is just the conditional expectation $\mathbb{E}[X\,|\,X^{\prime}=x^{\prime}]$ when $(X,X^{\prime})\sim\pi$. If the convex relaxation above is tight, then $\delta$ must be solution to (2.1). Otherwise, $\delta$ can be thought of as an approximate solution to (2.1).
2.3 Wasserstein Geometry
We next review some aspects of the geometry of the Wasserstein space; further information can be found in the same references as above, i.e., Villani (2009); Rachev and Rüschendorf (1998); Panaretos and Zemel (2020); Chewi et al. (2025).
We write $\mathcal{P}_{2}(\mathbb{R}^{k})$ for the space of probability measures $\mu$ on $\mathbb{R}^{k}$ satisfying $\int_{\mathbb{R}^{k}}\|x\|^{2}\textnormal{d}\mu(x)<\infty$. For any probability measures $\mu,\mu^{\prime}\in\mathcal{P}_{2}(\mathbb{R}^{k})$, we write $W_{2}^{2}(\mu,\mu^{\prime})$ for the optimal value of problem (2.2) above, which is called the (square of the) 2-Wasserstein distance between $\mu$ and $\mu^{\prime}$. (If $\mu$ additionally has a density with respect to Lebesgue measure, then we know that this is equal to the optimal value of problem (2.1).) The metric space $(\mathcal{P}_{2}(\mathbb{R}^{k}),W_{2})$ is called the 2-Wasserstein space.
These notions simplify significantly in the case of centered Gaussian measures, which will prove useful in our study of the variance-constrained denoising problem ($\mathcal{VCB}$). For $m\in\mathbb{N}$, we write $\mathbb{K}(m)$ for the space of (strictly) positive definite $m\times m$ matrices. For any $\Sigma,\Sigma^{\prime}\in\mathbb{K}(m)$, we may consider problem (2.1) for probability measures $\mu=\mathcal{N}(0,\Sigma)$ and $\mu^{\prime}=\mathcal{N}(0,\Sigma^{\prime})$; it turns out that this problem has a (necessarily unique) solution given by the linear map $\delta(x):=\mathbf{t}_{\Sigma}^{\Sigma^{\prime}}x$, where
$\mathbf{t}_{\Sigma}^{\Sigma^{\prime}}:=\Sigma^{-\nicefrac{{1}}{{2}}}(\Sigma^{\nicefrac{{1}}{{2}}}\Sigma^{\prime}\Sigma^{\nicefrac{{1}}{{2}}})^{\nicefrac{{1}}{{2}}}{\Sigma}^{-\nicefrac{{1}}{{2}}}.$ (2.3)
Here, we write $A^{\nicefrac{{1}}{{2}}}$ for the unique positive definite square root of a positive definite matrix $A\in\mathbb{K}(m)$. We also write $\preceq$ and $\prec$ for the usual Loewner order on $\mathbb{K}(m)$, that is, the order generated by the positive semi-definite (PSD) cone. The space $\mathbb{K}(m)$ endowed with the metric inherited from $W_{2}$, under the natural identification, is referred to as the Bures-Wasserstein space.

### v2

@@ Abstract
Keywords: constrained Bayes estimation; $G$-modeling; Gaussian mixture model; errors-in-variables regression; latent variable model; (smoothed) nonparametric maximum likelihood estimation; Wasserstein space
@@ 1.1 Problem Statement
More precisely, the goal of denoising is to construct a function $\delta:\mathbb{R}^{d}\to\mathbb{R}^{m}$ such that the risk $\mathbb{E}[\|\delta(Z)-\Theta\|^{2}]$ is small, with expectation taken over the joint distribution of $(\Theta,Z)$. If the distribution $G$ is known, then the best choice is the posterior mean $\delta_{\mathcal{B}}(z):=\mathbb{E}[\Theta\,|\,Z=z]$, called the Bayes denoiser (with respect to the squared error loss). If the distribution $G$ is not known, then it is often still possible to construct some function $\hat{\delta}_{\mathcal{B}}$, called an empirical Bayes (EB) denoiser (Robbins, 1956; Efron, 2019), which achieves a risk that is nearly equal to that of the Bayes denoiser.
@@
Similar considerations have been noted by many authors over the last hundred years, e.g., Eddington (1940); Louis (1984); Ghosh (1992); Loredo (2007); Bloom et al. (2017); Agarwal et al. (2020); Chen et al. (2025), and in textbooks [Raudenbush and Bryk, 2002, p. 88; Rao and Molina, 2015, p. 316].
@@ 1.2 Classical Denoising and Three Constrained Denoising Problems
We now describe the classical denoising problems and three constrained variants that are of interest in this work. First, we consider the (unconstrained) Bayes denoising problem,
$\underset{\delta:\mathbb{R}^{d}\to\mathbb{R}^{m}}{\textnormal{minimize}}\quad\mathbb{E}\left[\|\delta(Z)-\Theta\|^{2}\right],$ ($\mathcal{B}$)
whose solution is the posterior mean $\delta_{\mathcal{B}}(z)=\mathbb{E}[\Theta\,|\,Z=z]$, and we next add suitable constraints.
@@
$\underset{\delta:\mathbb{R}^{d}\to\mathbb{R}^{m}}{\textnormal{minimize}}\quad\mathbb{E}\left[\|\delta(Z)-\Theta\|^{2}\right]\qquad\textnormal{s.t.}\qquad\mathbb{E}[\delta(Z)]=\mathbb{E}[\Theta]\quad\textnormal{and}\quad\textnormal{Cov}(\delta(Z))=\textnormal{Cov}(\Theta).$ ($\mathcal{VCB}$)
Problem ($\mathcal{VCB}$) aims to find a denoiser $\delta$ that minimizes the Bayes risk subject to the constraint that the first two moments of $\delta$ match those of the latent variable $\Theta$, thus mitigating the shrinkage of (1.2).
@@
$\underset{\delta:\mathbb{R}^{d}\to\mathbb{R}^{m}}{\textnormal{minimize}}\quad\mathbb{E}\left[\|\delta(Z)-\Theta\|^{2}\right]\qquad\textnormal{s.t.}\qquad\delta(Z)\overset{\mathcal{D}}{=}\Theta.$ ($\mathcal{DCB}$)
@@
$\underset{\delta:\mathbb{R}^{d}\to\mathbb{R}^{m}}{\textnormal{minimize}}\quad\mathbb{E}\left[\|\delta(Z)-\Theta\|^{2}\right]\qquad\textnormal{s.t.}\qquad\mathbb{E}[\psi_{\ell}(\delta(Z))]=\mathbb{E}[\psi_{\ell}(\Theta)]\quad\textnormal{ for all }1\leq\ell\leq k$ ($\mathcal{GCB}$)
where $\psi_{1},\ldots,\psi_{k}:\mathbb{R}^{m}\to\mathbb{R}$ is some prescribed set of measurable functions. This formulation is general enough to cover a wide variety of problems; for example, it encompasses ($\mathcal{VCB}$) when $m=1$ by taking $\psi_{1}(z):=z$ and $\psi_{2}(z):=z^{2}$, it also encompasses ($\mathcal{DCB}$) in a limiting sense when $\psi_{1},\psi_{2},\ldots$ are taken to be a suitable sequence of bounded measurable functions. The added generality is useful in some novel applications and substantially extends the scope of constrained denoising.
@@ 1.3 Summary of Results
The main results of this paper concern the various constrained Bayes problems above, their EB approximations, extensions, and applications. We now describe these points in more detail.
@@
Distribution-Constrained Denoising. Next, we consider the distribution-constrained Bayes problem ($\mathcal{DCB}$). At the population level, we give conditions under which there is a uniquely defined distribution-constrained Bayes denoiser $\delta_{\mathcal{DCB}}$ (Theorem 3.8); at the empirical level, we provide an EB denoising scheme (Algorithm 2) which produces a denoiser $\hat{\delta}_{\mathcal{DCB}}$ that exhibits an explicit rate of convergence to the oracle constrained denoiser (Theorem 4.4). This work extends recent results in García-Trillos and Sen (2024) concerning the population-level problem; we provide rigorous statistical guarantees for an EB procedure which was stated (but not rigorously studied) in García-Trillos and Sen (2024, Appendix E). In Figure 1.1 (fourth column) we display the resulting denoisers in the simulation. The figure also highlights the difference between the problem of distribution-constrained EB denoising and the problem of deconvolution (Meister, 2009), that denoising implies a correspondence between the denoised data and the observations, while deconvolution does not.
@@
General-Constrained Denoising. Third, we consider the general-constrained Bayes problem ($\mathcal{GCB}$), which encapsulates a wide variety of practical constraints (for instance, the fourth-moment constraints of Armstrong et al. (2022) or support constraints such as nonnegativity). While the precise statements and discussions of these results can be found in Appendix C, we briefly summarize them here. At the population level, we give conditions under which there is a uniquely-defined general-constrained Bayes denoiser $\delta_{\mathcal{GCB}}$; at the empirical level, we provide an EB denoising scheme which produces a denoiser $\hat{\delta}_{\mathcal{GCB}}$ that exhibits an explicit rate of convergence to the oracle constrained denoiser.
@@
A fundamental flexibility of our paradigm is that it allows one to transform an arbitrary unconstrained EB denoiser into a suitable constrained EB denoiser, and the constrained denoiser inherits its rate of convergence from the unconstrained EB denoiser. For example, the unconstrained denoiser may arise via $G$-modeling, $F$-modeling, parametric EB, smoothed versions thereof, hybrids thereof, and more. This allows us to determine the rates of convergence for our constrained EB denoising procedures in many different settings of interest (e.g., light- and heavy-tailed $G$, Gaussian and Poisson likelihood $\{P_{\theta}\}_{\theta}$, conjugate parametric models), and we give a detailed analysis of the resulting rates of convergence in Appendix B. A notable example is the case of a nonparametric latent variable distribution and a Gaussian likelihood where we use nonparametric maximum likelihood (NPMLE) for $G$-modeling; our results imply that the rate of convergence of $\hat{\delta}_{\mathcal{VCB}}$ is nearly-parametric, and the rate of convergence of $\hat{\delta}_{\mathcal{DCB}}$ is slow (due to the slow minimax rate of nonparametric deconvolution, i.e., Carroll and Hall (1988); Zhang (1990); Fan (1991)).
@@
We also illustrate our methodology in two applications (Section 6). In an application to astronomy (Section 6.1), we use our results to denoise the distribution of latent chemical abundances in a large catalog of observed stars, furthering the analyses of Soloff et al. (2024); Ratcliffe et al. (2020). In an application to baseball (Section 6.2), we use our results to denoise the distribution of latent batting skill for players who are promoted from minor- to major-league play. Because these applications involve some degree of heterogeneity in the observations (e.g., heteroskedasticity in the Gaussian case) we include a detailed discussion (Section D) of some considerations for extensions to the heterogeneous case.
@@ 1.4 Related Literature
Our proposed methodology is closely tied to practical considerations in a number of application areas. In this paper, we specifically consider applications to astronomy (Section 6.1) and baseball (Section 6.2). Regarding astronomy, Loredo (2007) has argued that distributional constraints on denoising procedures are a fundamental feature of astrostatistics. Another application is to baseball statistics; while our specific application is novel, it complements a long tradition of EB methods in baseball (Efron and Morris, 1975; Brown, 2008; Jiang and Zhang, 2010; Gu and Koenker, 2017a). Our methodology is directly relevant to domains that have recognized the need for constrained denoising and have employed variance-constrained denoisers, including multisite experiments in education research (Raudenbush and Bloom, 2015; Bloom et al., 2017; Lee et al., 2024), small-area estimation (Devine et al., 1994; Ghosh and Rao, 1994; Leyland and Davies, 2005), epidemiology (Lyles et al., 1997; Moore et al., 2010), and errors-in-variables regression (Freedman et al., 2004). We also anticipate applications in other fields, such as those in labor economics reviewed by Walters (2024).
2 Preliminaries on Optimal Transport
@@
$\underset{T:\mathbb{R}^{k}\to\mathbb{R}^{k}}{\textnormal{minimize}}\int_{\mathbb{R}^{k}}\|x-T(x)\|^{2}\textnormal{d}\mu(x)\qquad\textnormal{s.t.}\qquad T_{\#}\mu=\mu^{\prime}.$ (2.1)
@@
$\underset{\pi\in\Gamma(\mu;\mu^{\prime})}{\textnormal{minimize}}\int_{\mathbb{R}^{k}}\|x-x^{\prime}\|^{2}\textnormal{d}\pi(x,x^{\prime})$ (2.2)
@@
In practice, a common approach to solving (2.1) is as follows. First, find a solution $\pi$ to (2.2). Second, define the function $\delta:\mathbb{R}^{k}\to\mathbb{R}^{k}$ via $\delta(x^{\prime}):=\int_{\mathbb{R}^{k}}x\,\pi(\textnormal{d}x,x^{\prime})\big{/}\int_{\mathbb{R}^{k}}\pi(\textnormal{d}x,x^{\prime})$ called the barycentric projection of $\pi$; note that this is just the conditional expectation $\mathbb{E}[X\,|\,X^{\prime}=x^{\prime}]$ when $(X,X^{\prime})\sim\pi$. If the convex relaxation above is tight, then $\delta$ must be solution to (2.1). Otherwise, $\delta$ can be thought of as an approximate solution to (2.1).
We next review some aspects of the geometry of the Wasserstein space. We write $\mathcal{P}_{2}(\mathbb{R}^{k})$ for the space of probability measures $\mu$ on $\mathbb{R}^{k}$ satisfying $\int_{\mathbb{R}^{k}}\|x\|^{2}\textnormal{d}\mu(x)<\infty$. For any probability measures $\mu,\mu^{\prime}\in\mathcal{P}_{2}(\mathbb{R}^{k})$, we write $W_{2}^{2}(\mu,\mu^{\prime})$ for the optimal value of problem (2.2) above, which is called the (square of the) 2-Wasserstein distance between $\mu$ and $\mu^{\prime}$. (If $\mu$ additionally has a density with respect to Lebesgue measure, then we know that this is equal to the optimal value of problem (2.1).) The metric space $(\mathcal{P}_{2}(\mathbb{R}^{k}),W_{2})$ is called the 2-Wasserstein space.
These notions simplify significantly in the case of centered Gaussian measures, which will prove useful in our study of the variance-constrained denoising problem ($\mathcal{VCB}$). For $m\in\mathbb{N}$, we write $\mathbb{K}(m)$ for the space of (strictly) positive definite $m\times m$ matrices. For any $\Sigma,\Sigma^{\prime}\in\mathbb{K}(m)$, we may consider problem (2.1) for probability measures $\mu=\mathcal{N}(0,\Sigma)$ and $\mu^{\prime}=\mathcal{N}(0,\Sigma^{\prime})$; it turns out that this problem has a (necessarily unique) solution given by the linear map $\delta(x):=\mathbf{t}_{\Sigma}^{\Sigma^{\prime}}x$, where $\mathbf{t}_{\Sigma}^{\Sigma^{\prime}}:=\Sigma^{-\nicefrac{{1}}{{2}}}(\Sigma^{\nicefrac{{1}}{{2}}}\Sigma^{\prime}\Sigma^{\nicefrac{{1}}{{2}}})^{\nicefrac{{1}}{{2}}}{\Sigma}^{-\nicefrac{{1}}{{2}}}$. Here, we write $A^{\nicefrac{{1}}{{2}}}$ for the unique positive definite square root of a positive definite matrix $A\in\mathbb{K}(m)$. We also write $\preceq$ and $\prec$ for the usual Loewner order on $\mathbb{K}(m)$, that is, the order generated by the positive semi-definite (PSD) cone. The space $\mathbb{K}(m)$ endowed with the metric inherited from $W_{2}$, under the natural identification, is referred to as the Bures-Wasserstein space.